6.00 metres of a wire yields a resistance of 84.0 ȍ. If the resistivity of the wire is 4.30 × 10–7 ȍm, find the diameter of the wire. 


1
Expert's answer
2021-12-23T10:50:17-0500

First, find the area:


"A=\frac{\rho L}{R}."

Second, find the diameter:


"D=\sqrt{\frac{4A}{π}}=2\sqrt{\frac{\rho L}{πR}}=4.4 \text{ mm}."